If the sides of a square are lengthened by 7 cm, the area becomes 144cm^2. Find the length of a side of the original square.
The length of a side of the original square is _____ cm.

Answer by rapaljer(4671)

Let x = original side
x+7 = increased side
Area = s^2 = (x+7)^2 = 144

Take the square root of each side:
x+7 = 12 (it can't be a negative 12!)
x=5 cm.

R^2